# UGround: Towards Unified Visual Grounding with Unrolled Transformers

### v1

@@ Appendix A Supplementary Statements
Figure 4: The UGround code repository.
The UGround code repository. As shown in Fig. 4, the UGround project is organized with a model directory for core implementations, a scripts directory containing training and testing scripts, and a configs directory for configuration files. The main model architecture is defined in UGround.py within the UGroundForCausalLM class, which extends LlavaLlamaForCausalLM and integrates a Segment Anything Model (SAM) for mask generation. A central component of this architecture is the PolicyPromptedMasking class, detailed in PPM.py, which is instantiated and used by the UGroundForCausalLM class to dynamically select hidden state features from different network layers. The selection policy is trained using the REINFORCE algorithm with either an EMA or a critic-based baseline. All codes and models are publicly available at https://github.com/rui-qian/UGround.
Figure 5: The UGround dashboard.
Visualization across layers 1-40. Fig. 68 present a qualitative analysis of similarity map activations across layers 1–40, illustrating their alignment with the ground-truth mask. Each figure contrasts heatmaps generated by prior works using a fixed last hidden layer (a) with those from our proposed Policy-Prompted Masking method (b). Unlike the baseline approach (a), our method introduces stochastic layer selection: in each forward pass, a layer $\ell^{*}$ is randomly chosen, and multiple $\mathcal{T}$ stochastic samples result in an approximate Bayesian uncertainty estimation, akin to Monte Carlo Dropout. This sampling mechanism explores diverse pathways, alleviates over-reliance on any single trajectory, and virtually forms an ensemble of subnetworks, thereby enhancing robustness. As a result, the similarity maps in (b) show activations that are sharply concentrated on the target object in intermediate and later layers, yielding more accurate alignment with the ground-truth mask compared to the diffuse and noise-prone activations observed in (a).

### v2

@@ Appendix F Ethics and Societal Impact
Appendix G Limitations and Future Work
We observe that UGround generally understands which object is being referred to, which means that the semantic grounding is correct (i.e., the language responses are accurate), yet the predicted masks often fail to capture the full spatial extent of the object. The masks are frequently fragmented, perforated with holes, spatially offset, or spilled into irrelevant regions; in some cases, they even collapse entirely. This indicates that the issue does not stem from semantic misinterpretation, but from imperfect cross-modal geometric alignment.
Although the <SEG> token is intended to represent the “target object,” it is fundamentally a language-space embedding whose feature distribution is not inherently aligned with the spatial structure of visual tokens. In other words, <SEG> token $\neq$ spatially aligned embedding, embedding space $\neq$ image space. For example, a distant patch of blue sky and a blue notebook on a desk may appear close in embedding space due to similar color or texture, even though they are far apart in the image. Thus, the <SEG> token captures semantic concepts rather than the spatial statistics of an object’s distribution. While using the mask as a prompt does introduce spatial cues, the internal decoding mechanism of SAM remains largely unexplored, making it unclear how these cues propagate into its actual segmentation behavior.
@@ Appendix G Limitations and Future Work
Figure 7: Visualization of UGround-7B on ReasonSeg test set. Zoom in for best view.
Figure 8: Visualization of UGround-13B on the ReasonSeg test set, demonstrating UGround’s capability in semantic segmentation, i.e., assigning pixel-level labels for a target semantic category (with background as 0). Zoom in for best view.
Figure 9: Visualization of UGround-7B on gRefCOCO (Liu et al., 2023) test sets A and B (I), demonstrating UGround’s capability in instance segmentation, i.e., predicting instance-level masks for individual object instances (with separate masks for objects of the same category). Zoom in for best view.
Figure 10: Visualization of UGround-7B on gRefCOCO (Liu et al., 2023) test sets A and B (II), demonstrating UGround’s capability in instance segmentation, i.e., predicting instance-level masks for individual object instances (with separate masks for objects of the same category). Zoom in for best view.
Figure 11: Visualization of UGround-7B on the FP-RefCOCO(+/g) val set, demonstrating UGround’s capability in handling false-premise queries, i.e., queries that refer to objects not present in the given image (empty target). Zoom in for best view.
Figure 12: The UGround dashboard, demonstrating a unified interface that supports multiple visual grounding models, i.e., LISA (Lai et al., 2024), SESAME (Wu et al., 2024), PixelLM (Ren et al., 2024), GSVA (Xia et al., 2024), and READ (Qian et al., 2025). This example illustrates the result of a single UGround inference.
Figure 13: The UGround code repository. All codes and models are publicly available at https://github.com/rui-qian/UGround.